OpenStudy (anonymous):

Write your answer in this form, without leaving spaces: (2x-2)(x+1). Factor the trinomial: 3x^2+11x+6

OpenStudy (cwrw238):

(3x + 2)(x + 3)

OpenStudy (anonymous):

\[3x^{2}+11x=6\] \[3x^{2}+9x+2x+6\] \[3x\left( x+3 \right)+2\left( x+3 \right)\] \[\left( 3x+2 \right)\left( x+3 \right)\]